In each reaction box, place the best reagent or reactant from the list below. Reagents may be used more than once or not at all. Draw the intermediate products B and C (both are neutral; omit byproducts).

media/b56/b569d924-e14c-485f-a78d-e0
Answer #1
Concepts and reason

One or more substances that react together to form a product is called a chemical reaction. Most chemical reactions occurs with the help of reagents.

Reagents are the substances used to transform one form of substance to another. The reaction undergoes series of steps

Deprotonation of alkyne, substitution and hydrogenation reaction to attain a final product.

Fundamentals

The reagents play a vital role in a reaction to occur. Each reagent plays a different role in the chemical reaction.

Deprotonation of alkynes: Alkyne undergoes Deprotonation with the help of a reagent sodium amide (NaNH2)\left( {{\rm{NaN}}{{\rm{H}}_{\rm{2}}}} \right) . Acetylide ion forms when Sodium amide is reacted with alkyne. Acetylide ion acts as good nucleophile in which it undergoes substitution reaction with alkyl halides and forms a new carbon chain.

Substitution: The Nucleophilic substitution reaction (SN2)\left( {{{\rm{S}}_{\rm{N}}}{\rm{2}}} \right) occurs when the nucleophile undergoes a reaction with alkyl halides.

Hydrogenation: The addition of H2{{\rm{H}}_{\rm{2}}} in the presence of palladium is used to reduce alkenes and alkynes to alkanes.

NaNH
HC=co Nat
H-C=—H
A acetylene

اند که
HẨc=C: Na
Br
د
-H C=
C
1-bromo-2-methylpropane
B

NaNH2
Hz,
Sac
H-CEC-

Galetal
Na:
CC-
2 - butanone

H, Pd/C

Ans:

A
acetylene
1.1) NINH,1
1.2) 1-bromo-2-methylpropane
H-CEC-
2.1)
NaNH,
2.2)
2-butanone
step 3 reagent
H, Pd/C
3,7-dimethyloct